a) Given the following simple Keynesian Model: Y = C + I + G + X-M, where Consumption schedule is given as C= 100 +0.75Y Investment (I) = 50 Government (G) = 100 and Net Export (X-M) = 20 i. Calculate the Equlibrium Level of Income [4 Marks] ii. Calculate the size of Consumption at the Equilibrium Level [2 Marks] iii. Calculate the value of the Government Multiplier [2 Marks] iv. Assume Investment (I) changes by 50; calculate the new equilibrium level of Income


1
Expert's answer
2021-05-17T10:38:26-0400

(i)

 "Y = C + I + G + X-M"

"Y = 100+0.75Y+50+100+20"

"Y = 270+0.75Y"

"0.25Y = 270"

"Y = 1080"

Equilibrium level of Income is 1080


(ii)

"C = 100+0.75Y"

"C = 100+0.75 \times 1080"

"C = 100+810"

"C = 910"

Size of consumption at equilibrium level is 910


3.

Government Multiplier "= \frac{1}{(1-MPC) }= \frac{1}{(1-0.75)} = \1/0.25 = 4"

Value of Government multiplier is 4


4.

I = 100

Therefore, 

"Y = 100+0.75Y+100+100+20"

"Y = 320+0.75Y"

"0.25Y = 320"

"Y = 1280"

Therefore, new level of equilibrium income is 1280